To buy both a new car and a new house, Tina sought two loans totalling $332,888. The simple interest rate on the first loan was 8.9%, while the simple interest rate on the second loan was 6.0%. At the end of the first year, Tina paid a combined interest payment of $20,230.51. What were the amounts of the two loans?
Answer by rfer(16322)

0.089x+0.06(332888-x)=20230.51
0.089x+19973.28-0.06x=20230.51
0.029x=257.23
x=257.23/0.029
x=8870 at 8.9%
332888-8870=324018 at 6%